OpenStudy (anonymous):

A circle has a diameter with end points (3,-2) and (9,-8) what is the equation of the circle? My options: R^2=(x+6)^2+(y-5)^2 R^2=(x+6)^2+(y-3)^2 R^2=(x-6)^2+(y+3)^2 R^2=(x-6)^2+(y+5)^2

OpenStudy (mertsj):

Find the midpoint of the diameter. That will be the center of the circle.

OpenStudy (anonymous):

Equation of any circle centred on the point (a,b) with a radius r is given by:\[(x-a)^2 + (y-b)^2 = r^2\] So in your case, the radius is just R (still unknown). And we are left to find the centre ourselves.  We see that the centre must be the middle of this line one the diagram. So taking the average of the x and y values gives us (9+3)/2=6, and (-2-8)/2=-5. So the centre is (6,-5). From here plug it back into the formula and we get the formula to be \[(x-6)^2 + (y+5)^2 = R^2\]

OpenStudy (anonymous):

Thank you!